OpenStudy (anonymous):

y+x^2=2x y+2x=4 find the solutions

myininaya (myininaya):

Solve the bottom one for 2x by subtracting y on both sides

myininaya (myininaya):

Then you have both equations =2x

OpenStudy (anonymous):

y=4-2x put it into y+x^2=2x (4-2x)+x^2=2x x^2-2x+4=2x x^2+4=0

myininaya (myininaya):

So that means whatever they equal must equal

myininaya (myininaya):

Or solve for y like cyclets did

OpenStudy (anonymous):

thanks

myininaya (myininaya):

And also one thing there is an error -2x-2x=-4x

OpenStudy (anonymous):

ops..sorry.. y=4-2x put it into y+x^2=2x (4-2x)+x^2=2x x^2-2x+4=2x x^2-4x+4=0 (x+2)(x-2)=0

OpenStudy (anonymous):

(x-2)(x-2)=0

OpenStudy (anonymous):

so would the points be 2,0 and -2,8

OpenStudy (anonymous):

the point is 0,0 to 2,0

OpenStudy (anonymous):

wouldnt the x's be 2 and -2 though

myininaya (myininaya):

I think he meant (x-2)(x-2)=0 which will get you one zero with multiplicity 2 x=2 So to find what y is just plug in 2 for x and solve for y (just use one of the equations given)

OpenStudy (anonymous):

yes..thanks myininaya ~ appreciate that